OpenStudy (anonymous):

Determine whether the point (1,5) is a solution to the system of inequalities below y>3x Y>2x+1

OpenStudy (anonymous):



OpenStudy (anonymous):

thats for the second one

OpenStudy (anonymous):

9x^4.

OpenStudy (anonymous):

plug in 5 for y and 1 for x and see if the inequalities are true and they are

OpenStudy (anonymous):

4x + 7 = 9x - 4 9x - 4x = 7 + 4 5x = 11 x = 11/5 or 2 1/5 Answer: x = 11/5 or 2 1/5 Proof: 4(11/5) + 7 = 9(11/5) - 4 44/5 + 35/5 = 99/5 - 20/5 79/5 = 79/5

OpenStudy (lgbasallote):

you guys do know you're in the writing section right? lol